# ReSAM: Refine, Requery, and Reinforce: Self-Prompting Point-Supervised Segmentation for Remote Sensing Images

### v1

@@ 3.4 Requery: Self-Prompting with Box Generation
We employ a self-prompting mechanism that automatically re-queries SAM using the pseudo masks generated in the refinement stage. For each instance $\tilde{M}_{ij}^{(k)}$, we compute its minimal enclosing bounding box:
$B=\text{Box}(\tilde{M}_{ij}^{(k)}),$ (11)
@@
$M_{r}=\Phi_{\text{mask}}(\Phi_{\text{img}}(I_{w}),\Phi_{\text{prompt}}(P_{B})).$ (12)
This self-prompting step effectively converts uncertain point supervision and removes noise in pseudo-label generation into structured region queries, producing refined and context-aware masks that act as pseudo ground truth for training.
@@ 3.5 Reinforce: Soft Semantic Alignment
Given the embeddings of objects from the current iteration, we first stack them into a memory queue $q$ containing embeddings from recent iterations. Let $f_{i}$ denote the $i$-th embedding in the queue. We then compute pairwise cosine similarities among all embeddings in the queue:
@@
$\mathcal{L}_{\text{SSAL}}=\frac{1}{N}\sum_{i}\sum{j\neq i}(1-S_{ij}),p_{ij}.$ (15)
@@
The memory queue not only stabilizes training but also enables the model to consolidate semantic prototypes over time, ensuring robust learning even from imperfect pseudo-labels.
@@ 3.6 Total Objective
We optimize a composite loss integrating segmentation and semantic stability, overall, we minimize:
@@
where $\mathcal{L}_{\text{focal}}$, $\mathcal{L}_{\text{dice}}$, and $\mathcal{L}_{\text{iou}}$ supervise pixel-wise mask quality, and $\mathcal{L}_{\text{SSAL}}$ enforces cross-iteration feature consistency.
@@ 4.1 Datasets
We evaluate our proposed method on multiple instance level remote sensing benchmark datasets using minimal point supervision. Our experiments demonstrate that self-prompting significantly improves the segmentation accuracy for SAM. We conduct experiments on three different datasets:
  • NWPU VHR-10: dataset  [cheng2016learning] consists of 800 very-high resolution optical images (10 Object classes). Following prior work, we treat the problem as class-agnostic instance segmentation and use 520 images for training and 130 for testing which are a positive subset of this dataset.
@@
  • WHU: dataset  [ji2018fully] is a large-scale building segmentation dataset with high spatial resolution (0.075m). We used the standard training/validation split provided with the dataset, which splits the data into 4736 for training and 1036 for testing.
@@ 4.2 Experimental Settings
For all experiments, we use sparse point supervision only. We randomly sample positive and negative points from GT mask for each instance (same as our baseline [liu2025pointsam]). We experiment with 1,2, and 3 points sampled per instance. No annotated full mask or box was provided to the model during training.
@@
We compare our method to several baselines for weakly supervised and recent point-based segmentation for SAM. We compare our method with Direct test a pretrained SAM without adaptation, prompted with ground-truth points. Self-Training [liu2102unbiased] uses student-teacher structure without any regularizers. WeSAM [zhang2024improving], DePT [gao2022visual] and Tribe [su2024towards] representative source-free domain adaptation utilized prompt-based direction feeding during training without source data. PointSAM [liu2025pointsam] employs Prototype-based self-training method that uses FINCH[sarfraz2019efficient] clustering and negative prompt calibration. Additionally, we also report the Supervised upper bound, Full-mask finetuning (LoRA) to indicate the upper limit.
@@
We fine-tune the LoRA module of SAM image encoder with rank=4. Only LoRA parameters are updated with Adam optimizer across all experiments. To validate our method with limited resources, training is performed with learning rate of $5\times 10^{-4}$, weight decay $1\times 10^{-4}$ on RTX 3060 with batch size of 1. We set $\alpha$ and $\beta$ in Eq.16 to 20 and 0.1 for focal loss and SSA loss, respectively. An instance embedding queue $q$ (FIFO) of length 32 is maintained. Embedding features are L2-normalized before being pushed the o queue. From augmented strategy in [zhang2024improving] we apply both strong and weak augmentation for better domain adaptation. To maintain the soft alignment of top instance object embedding, we set $\tau$ to 0.05. For entropy threshold $\epsilon$, we set it to 0.3.
@@ 4.3 Quantitative Results
Tables 24 compare ReSAM with state-of-the-art prompt-based segmentation methods on NWPU VHR-10, WHU, and HRSID-Inshore. ReSAM consistently outperforms SAM- and SAM2-based baselines, demonstrating strong generalization and effective pseudo-label adaptation. On NWPU VHR-10, it surpasses all competitors, with gains up to +2.0 IoU and +1.8 F1 over PointSAM, and performance improves steadily from 1- to 3-point inputs, narrowing the gap with fully supervised models to under 12 IoU points. On WHU, ReSAM achieves top results in most settings, particularly under 2-point prompts, though 3-point SAM2 slightly reduces precision due to redundant spatial priors. On HRSID-Inshore, single-point performance is highest, but multi-point prompts show instability likely caused by cluttered backgrounds and small objects. Overall, ReSAM delivers consistent gains across datasets, validating its refined pseudo-label generation and re-prompting strategy.
@@ 5 Ablation Studies
We perform an ablation to isolate the influence of each component of ReSAM.
@@
The ablation results in Table 5 show that each component of our framework contributes to suppressing overlapping or ambiguous masks. Self-training provides an initial improvement over Direct SAM, but adding Requery yields a larger gain by repeatedly resolving uncertain regions and reducing boundary conflicts. SSA offers a complementary boost by enforcing softer semantic consistency in feature space, which helps prevent redundant or overlapping predictions. Fine-tuning with LoRA further stabilizes these corrections, outperforming Adapter and LayerNorm-based alternatives. The full ReSAM model, which combines Requery, SSA, and LoRA, achieves the highest mIoU (73.4%, +13.4) on WHU dataset, especially for 1 point prompts, demonstrating that spatial requerying, semantic alignment, and efficient parameter tuning jointly produce stronger overlap suppression and more accurate final masks.
@@
Full ReSAM (Requery + SSA + LoRA) 73.4 +13.4
@@
To evaluate the effectiveness of the proposed Semantic Spatial Alignment (SSA) module, we conduct an ablation study on the HRSID and WHU datasets across the first five training epochs. As shown in Fig. 5, integrating SSA consistently improves segmentation performance on both benchmarks. On HRSID, the model with SSA achieves higher mIoU at every epoch, increasing from 46.8% to 58.4%, whereas the baseline without SSA plateaus around 55% and even declines after epoch 3. This demonstrates that SSA effectively mitigates feature misalignment in high-resolution scenes, leading to more stable optimization. A similar trend is observed on WHU: adding SSA boosts performance from 61.0% to 73.4%, outperforming the w/o SSA setting which saturates near 69.4%. The sharper improvement on WHU indicates that SSA particularly benefits datasets with large-scale structural variations by enforcing better spatial–semantic consistency. Overall, these results confirm that SSA provides a robust enhancement to feature alignment and yields consistent gains in segmentation accuracy across datasets.
@@ 6 Conclusion
We presented ReSAM, a self-prompting, point-supervised framework that converts sparse point annotations into high-quality box prompts through an iterative Refine–Requery–Reinforce (R³) loop. Coupled with Soft Semantic Alignment (SSA), ReSAM ensures semantic consistency while avoiding dense supervision and heavy memory costs. Experiments on WHU, HRSID, and NWPU VHR-10 demonstrate robust adaptation and improved segmentation over vanilla SAM and prior point-supervised methods. ReSAM relies on careful tuning of the iterative loop and may be less effective in extremely cluttered scenes or for objects with highly irregular shapes.

### v2

@@ 3.4 Requery: Self-Prompting with Box Generation
We employ a self-prompting mechanism that automatically requeries SAM using the pseudo masks generated during the refinement stage. For each instance $M^{\text{ref},(k)}_{ij}$, we compute its minimal enclosing bounding box:
$B=\text{Box}(M^{\text{ref},(k)}_{ij})$, (11)
@@
$M_{p}=\Phi_{\text{m}}(\Phi_{\text{i}}(I_{w}),\Phi_{\text{p}}(P_{B})).$ (12)
This self-prompting step effectively transforms uncertain point supervision and removes noise in pseudo-label generation into structured region queries, producing refined and context-aware masks that act as pseudo ground truth $M_{p}$ for training.
@@ 3.5 Reinforce: Soft Semantic Alignment
Given the embeddings of objects from the current iteration, we first stack them into a memory queue $q$ for each box prompts that contains embeddings from recent iterations. Let $f_{i}$ denote the $i$-th embedding in the queue. We then compute pairwise cosine similarities among all embeddings in the queue:
@@
$\mathcal{L}_{\text{SSAL}}=\frac{1}{N}\sum_{i}\sum_{j\neq i}(1-S_{ij})p_{ij}$ (15)
@@
The memory queue not only stabilizes training, but also enables the model to consolidate semantic prototypes over time, ensuring robust learning even from imperfect pseudo-labels.
@@ 3.6 Total Objective
We optimize a composite loss that integrates segmentation and semantic stability. Overall, we minimize
@@
where $\mathcal{L}_{\text{focal}}$, $\mathcal{L}_{\text{dice}}$, and $\mathcal{L}_{\text{iou}}$ supervise pixel-wise mask quality, and $\mathcal{L}_{\text{SSAL}}$ enforces cross-iteration feature consistency. We set the values of $\alpha$ and $\beta$ to 20 and 0.1.
@@ 4.1 Datasets
We evaluate our proposed method on multiple instance level remote sensing benchmark datasets using minimal point supervision. Our experiments demonstrate that self-prompting significantly improves the segmentation accuracy of SAM. We conduct experiments on three different datasets:
  • NWPU VHR-10: dataset  [cheng2016learning] consists of 800 very-high-resolution optical images (10 objects classes). Following prior work, we treat the problem as class-agnostic instance segmentation and use 520 images for training and 130 for testing which constitute a positive subset of this dataset.
@@
  • WHU: dataset  [ji2018fully] is a large-scale building segmentation dataset with high spatial resolution (0.075m). We used the standard training/validation split provided with the dataset, which includes 4736 images for training and 1036 for testing.
@@ 4.2 Experimental Settings
For all experiments, we use sparse point supervision only. We randomly sample positive and negative points from the ground truth (GT) mask for each instance (same as our baseline [liu2025pointsam]). We experiment with 1, 2, and 3 points sampled per instance. An annotated full mask or box was not provided to the model during training.
@@
We compare our method to several baselines for weakly supervised and recent point-based segmentation methods based on SAM. We compare our method with Direct test a pretrained SAM without adaptation, prompted with ground-truth points. Self-Training [liu2102unbiased] uses student-teacher structure without any regularizers. WeSAM [zhang2024improving], DePT [gao2022visual] and Tribe [su2024towards] representative source-free domain adaptation methods that utilized prompt-based direction feeding during training without source data. PointSAM [liu2025pointsam] employs Prototype-based self-training method that uses FINCH[sarfraz2019efficient] clustering and negative prompt calibration. Additionally, we also report the Supervised upper bound of Full-mask finetuning (LoRA) to indicate the upper limit.
@@
We fine-tune the LoRA module of SAM image encoder with rank=4. Only LoRA parameters were updated with Adam optimizer across all experiments. To validate our method with limited resources, training is performed with learning rate of $5\times 10^{-4}$, weight decay $1\times 10^{-4}$ on RTX 3060 with a batch size of 1. An instance embedding queue $q$ (FIFO) of length 32 is maintained. Embedding features are L2-normalized before being pushed to the queue. Following the augmented strategy in [zhang2024improving] we apply both strong and weak augmentations for better domain adaptation. To maintain the soft alignment of top instance object embedding, we set $\tau$ to 0.05.
@@ 4.3 Quantitative Results
Tables 24 compare ReSAM with state-of-the-art prompt-based segmentation methods on the NWPU VHR-10, WHU, and HRSID-Inshore datasets. ReSAM consistently outperforms SAM- and SAM2-based baselines, demonstrating strong generalization and effective pseudo-label adaptation. On NWPU VHR-10, it surpasses all competitors, with gains up to +2.0 IoU and +1.8 F1 over PointSAM, and performance improves steadily from 1- to 3-point inputs, narrowing the gap with fully supervised models to under 12 IoU points. On WHU, ReSAM achieves top results in most settings, particularly under 2-point prompts, though 3-point SAM2 slightly reduces precision due to redundant spatial priors. On HRSID-Inshore, single-point performance is highest, but multi-point prompts show instability, which is likely caused by cluttered backgrounds and small objects. Overall, ReSAM delivers consistent gains across datasets, validating its refined pseudo-label generation and re-prompting strategy.
@@ 5 Ablation Studies
The ablation results in Table 5 show that each component of our framework contributes to suppressing overlapping or ambiguous masks. Self-training provides an initial improvement over the Direct SAM, but adding Requery yields a larger gain by repeatedly resolving uncertain regions and reducing boundary conflicts. SSA offers a complementary boost by enforcing softer semantic consistency in feature space, which helps prevent redundant predictions. Fine-tuning with LoRA further stabilizes these corrections, outperforming Adapter and LayerNorm-based alternatives. The full ReSAM model, which combines Requery, SSA, and LoRA, achieves the highest mIoU (73.4%, +13.4) on WHU dataset, especially for 1 point prompts, demonstrating that spatial requerying, semantic alignment, and efficient parameter tuning jointly produce stronger overlap suppression and more stable final masks.
@@
Full ReSAM (R³) 73.4 +13.4
@@
To evaluate the effectiveness of the proposed Semantic Spatial Alignment (SSA) module, we performed an ablation study on the HRSID and WHU datasets over the first five training epochs. As shown in Fig. 5, integrating SSA consistently improves segmentation performance on both benchmarks. On HRSID, the model with SSA achieves higher mIoU at every epoch, increasing from 46.8% to 58.4%, whereas the baseline without SSA plateaus around 55% and even declines after epoch 3. This demonstrates that SSA effectively mitigates feature misalignment, leading to more stable optimization. A similar trend is observed on WHU: adding SSA boosts performance from 61.0% to 73.4%, outperforming the w/o SSA setting which saturates near 69.4%. The sharper improvement on WHU indicates that SSA particularly benefits datasets with large scale structural variations by enforcing better spatial–semantic consistency. Overall, these results confirm that SSA provides a robust enhancement to feature alignment and yields consistent gains in segmentation accuracy across datasets.
@@ 6 Conclusion
We present ReSAM, a self-prompting, point-supervised framework that converts sparse point annotations into high quality box prompts through an iterative Refine–Requery–Reinforce (R³) loop. Coupled with Soft Semantic Alignment (SSA), ReSAM ensures semantic consistency while avoiding dense supervision and heavy memory costs. Experiments on WHU, HRSID, and NWPU VHR-10 demonstrate robust adaptation and improved segmentation over vanilla SAM and prior point supervised methods.
Limitation and Future work
ReSAM performs best when the iterative loop is well tuned, but its effectiveness can decrease for irregularly shaped objects, particularly if the baseline model already struggles on the target dataset. Additionally, our method faces challenges in the 3-point adaptation, where performance degrades noticeably. We believe this is linked to SAM’s inherent limitation when dealing with densely distributed objects, which restricts the effectiveness of sparse point supervision and needed to investigate further.